OpenStudy (anonymous):

The system graphed is y is less than or equal to 5x+1 and x is less than or equal to 0 Which of the ordered pairs is a solution of the system? graph of the system y≤5x+1 and x≤0 A. (0, -1) B. (-2, 3) C. (2, 4) D. (3, -1)

OpenStudy (callisto):

consider x≤0, when x>0, the answer can be rejected therefore C and D will not be the answer consider y≤5x+1 when x = 0 RHS = 5(0)+1 = 1 > -1 therefore A can be the order pair when x=-2 RHS = 5(-2)+1 = -10+1 = -9<-1 therefore B does not satisfy the inequality So you should get the answer now